# DIRAS: Efficient LLM-Assisted Annotation of Document Relevance in Retrieval Augmented Generation

### v1

@@ 3.2 Open-Sourced LLM Fine-Tuning
Small-embed - - 83.63 66.34 -
Large-embed - - 84.07 69.36 -
BGE-Gemma - - 78.21 68.47 -
GPT-3.5 29.71 45.27 87.79 74.16 59.23
GPT-4 54.01 86.32 94.41 88.48 80.80
Llama3-CoT-Ask 36.57 76.58 93.32 86.15 73.16
Llama3-CoT-Tok 41.74 76.58 90.82 85.96 73.78
Llama3-Ask 40.18 82.11 94.04 86.02 75.59
Llama3-Tok 41.60 82.11 94.41 89.19 76.83
Phi3-CoT-Ask 36.08 72.95 92.35 80.56 70.48
Phi3-CoT-Tok 35.49 72.95 88.15 80.64 69.31
Phi3-Ask 32.30 73.23 91.16 80.05 69.18
Phi3-Tok 38.00 73.23 92.05 86.94 72.55
Gemma-CoT-Ask 31.60 72.38 91.14 81.39 69.13
Gemma-CoT-Tok 39.03 72.38 88.58 80.33 70.08
Gemma-Ask 25.74 67.13 88.27 77.43 64.64
Gemma-Tok 50.72 67.13 92.44 81.17 72.87
@@
Table 2: Comparison between the fine-tuned $\mathcal{M}_{o}$ and different baselines on ChatReport test data. Unc. denotes the Average Precision (AP) of predicting uncertainty; Bin. denotes F1 score of binary relevance prediction; Cal. is the average of AUROC, ECE, and Brier Score; Info. is the average of nDCG and MAP; Avg. denotes the average of all metrics. The best scores are bolded and the second bests are underlined.$\dagger$ denotes the best score achieved by each LLM architecture.
@@
DIRAS data fine-tuned models $\mathcal{M}_{o}$ will be used to predict all (query, document) combinations to mitigate annotation selection bias (Thakur et al., 2021). Therefore, shorter generations from $\mathcal{M}_{o}$ (e.g., without CoT) are favored as the inference cost for transformers increases quadratically with generation length. Additionally, the choice of calibration method matters for LLMs (Tian et al., 2023). To explore these aspects, we fine-tune $\mathcal{M}_{o}$ in four settings: $\mathcal{M}_{o}$-CoT-Ask, $\mathcal{M}_{o}$-CoT-Tok, $\mathcal{M}_{o}$-Ask, $\mathcal{M}_{o}$-Tok, where CoT means $\mathcal{M}_{o}$ is tuned to generate [Reason], [Guess], and [Confidence]; without CoT denotes $\mathcal{M}_{o}$ is tuned to only generate [Guess] and [Confidence]; “Ask” means the result is calibrated by the generated confidence score in [Confidence] field; and “Tok” means we take the token-level probability of “Yes/No” after “[Guess]:” as the confidence score for calibration. The prompt in Fig. 2 is used for fine-tuning. The “[Reason]:” line is removed in settings without CoT.
@@
We fine-tune Llama-3-8B-instruct (AI@Meta, 2024), gemma-7b-it (Team et al., 2024b), and Phi-3-mini-4k-instruct (Abdin et al., 2024) (details in App. F). We compare these fine-tuned models with baselines including GPT-3.5 and GPT-4 using prompt $\mathcal{P}$; the OpenAI embedding models text-embedding-3-small, and text-embedding-3-large; and BGE Gemma reranker [^5] https://huggingface.co/BAAI/bge-reranker-v2-gemma, a popular LLM-based reranker for general domain. As Fig. 3 shows, fine-tuning improves original models in all settings. Furthermore, Table 2 shows the results of all fine-tuned models in comparison to all baselines. We observe that $\mathcal{M}_{o}$-Tok outperforms other settings for all LLM architectures. The best setting Llama3-Tok achieves GPT-4 level performance in calibration and IR on unseen questions and reports.
@@
Interestingly, we find that omitting the chain of thought usually leads to a performance increase for all three LLM architectures. CoT sometimes leads to a limited increase when asking for calibration (Ask), but constantly results in a performance drop when calibrated with token-level probability (Tok). Therefore, $\mathcal{M}_{o}$ should be fine-tuned without CoT for performance and inference efficiency. Moreover, Tok rarely underperforms Ask, different from Tian et al. (2023)’s finding and our observations in Table 1. Thus, future work may consider probabilities of important tokens (e.g., Yes/No in our prompt template) as a promising calibration tool.
@@
4 Experiments on ClimRetrieve
In this section, we showcase how the DIRAS fine-tuned $\mathcal{M}_{o}$ can assist the IR annotation in a real-world setting, leveraging the ClimRetrieve dataset (Schimanski et al., 2024b). This dataset records analysts’ real-life procedure of sustainability report analyses: raise questions about sustainability reports and then go through these reports to find relevant information to answer their questions. The dataset comprises 43k (query, document) pairs, among which 595 pairs are annotated as relevant. Relevant documents are annotated with a relevance score from 1 to 3 (1 translates to partially relevant and 2/3 means relevant). App. I gives a detailed overview of the dataset.
Compared to ChatReport data, ClimRetrieve is a more challenging and realistic setting because: (1) For ChatReport data, we draft explicit relevance definitions and annotate relevance dependent on them. But for ClimRetrieve, the analysts’ mental model for what is (ir)relevant for their posed questions is unknown to us. (2) ClimRetrieve records human analysts’ real-life workflow of reading the full reports and searching for relevant information, which differs from ChatReport data where annotators are presented with (query, document) pairs as separate data points. Thus, ClimRetrieve only has gold labels for relevant (query, document) pairs. Other not annotated (query, document) combinations might be either irrelevant or a part of annotation selection bias – a widely existing problem in IR annotation (Thakur et al., 2021).
This challenging and realistic setting of ClimRetrieve allows us to investigate the following research questions regarding DIRAS fine-tuned $\mathcal{M}_{o}$: (1) RQ1: Can $\mathcal{M}_{o}$ understand fine-grained differences in degree of relevance? (2) RQ2: Can we improve $\mathcal{M}_{o}$’s performance through improving relevance definitions? (3) RQ3: Can $\mathcal{M}_{o}$ assists in mitigating annotation selection bias (Thakur et al., 2021)? (4) RQ4: Can $\mathcal{M}_{o}$’s predictions help benchmarking IR algorithms? We use the best $\mathcal{M}_{o}$ in § 3 (Llama-3 without CoT) to study these RQs.
4.1 RQ1: Understanding Fine-Grained Relevance Levels
@@
We first evaluate fine-tuned Llama-3 on 595 gold labels of ClimRetrieve to verify whether it can effectively recover analysts’ ranking for relevant content by understanding which documents are more helpful than others. Relevance definitions are drafted with GPT-4 with the same procedure as § 3. We report nDCG [^6] MAP can only measure binary relevance and since we only investigate relevant samples, it cannot be calculated. scores to measure the ranking performance on ClimRetrieve. Gold labels 1, 2, and 3 are assigned with relevance scores 1/3, 2/3, and 1. Besides OpenAI 3rd generation embedding models, we also have a random baseline where all (query, document) pairs are assigned a random relevance score between 0 and 1. The random baseline results are averaged over 5 random seeds (40 to 44). Importantly, all ClimRetrieve annotations are to some degree relevant, so improvement over the random baseline is challenging as the system needs to understand the trivial different degrees of relevance.
Table 3 presents different systems’ performance. There is a clear trend of outperformance of the fine-tuned Llama-3 models in this challenging setting. They also exceed the random baseline by a significant margin, indicating the model correctly understands the fine-grained levels of relevance.
4.2 RQ2: Improving Performance through Improving Definitions
@@
We then test whether we can improve $\mathcal{M}_{o}$’s performance by improving the relevance definition presented in prompts. In § 3 and § 4.1, we use GPT-4 to draft the relevance definitions. To investigate the role of the definitions, we compare two setups: (1) The generic relevance definition: the definition drafted by GPT-4. (2) The informed relevance definition: we use the labeled texts in the ClimRetrieve dataset as examples to create artificial expert- textitinformed relevance definitions. This way, we simulate the inclusion of expert knowledge to improve the relevance definition (see App. J for details).
@@
After creating the definitions, we repeat predicting the relevance scores with the DIRAS fine-tuned Llama-3. First, we repeat the setup of previous § 4.1 with only human-annotated 595 relevant (query, document) pairs. We find the utilization of expert-informed definitions produces similar results, at most improving nDCG (see App. K). We argue that the definition might only make a significant difference when predicting all (query, document) pairs instead of only the human-annotated ones. The inclusion of specific nuances might especially help when compared with non-relevant documents. Calculating the nDCG and MAP score for all 43k (query, document) pairs, we find evidence for this notion (see Table 4). Thus, the inclusion of expert-informed definitions seems to improve the performance (for more details, see App. K).
4.3 RQ3: Mitigating Annotation Bias
ClimRetrieve employs a real-world analyst scenario. This entails that the human only selectively annotates documents that are likely to be relevant and assumes unannotated documents as irrelevant (see e.g., Thakur et al., 2021). Therefore, the dataset allows us to investigate our model’s capabilities to counteract biases. For this purpose, we reannotated 288 strategically sampled (sampling details in App. M) (document, query) pairs to investigate Llama3-Tok’s performance on ClimRetrieve’s missing annotation. For samples not labeled as relevant in ClimRetrieve but by Llama3-Tok, the reannotations indicate that we can effectively overrule these blind spots. As Fig. 4 shows, increasing the threshold of Llama3-Tok confidence allows us to be increasingly sure that we indeed find a relevant document. For the around 60% of the samples with a Llama3-Tok confidence higher than 99%, almost all samples are reannotated as relevant. Besides annotation selection bias, there are also other reasons for overruling the initial annotations. ClimRetrieve annotators may rather have focused on hard relevant documents omitting partially relevant ones. Furthermore, our GPT-4 created definitions may broaden the scope of what is relevant. However, we find that our model is well-calibrated and consistent with the provided relevance definitions. Therefore, Llama3-Tok is helpful for mitigating annotation selection bias (for a detailed analysis, see App. M).
Figure 4: This table shows the percentage of agreement between Llama3-Tok and our human reannotation, and the amount of data remaining, when adjusting the confidence threshold for (query, document) pairs that were considered relevant by Llama3-Tok but not by the ClimRetrieve annotation.
4.4 RQ4: Using $\mathcal{M}_{o}$ to Benchmark IR
@@
Table 5: Different embedding models’ performance benchmarked by $\mathcal{M}_{o}$’s prediction on all 43K (query, document) pairs of ClimRetrieve. “ft” denotes the model is fine-tuned on in-domain data.
@@
After validating $\mathcal{M}_{o}$’s performance, we can use it to annotate all 43k ClimRetrieve datapoints and obtain a benchmarking dataset to select IR algorithms. This approach can be especially helpful when lacking human annotation and annotation selection bias is prevalent. Specifically, we compare the performance of embedding models before and after in-domain fine-tuning. If the $\mathcal{M}_{o}$-annotated benchmark gives higher scores to the fine-tuned checkpoints, that means it is capable of selecting a better model for this specific domain.
@@
For this experiment, we first fine-tune open-sourced embedding models bge-large-en-v1.5 and bge-base-en-v1.5 (Chen et al., 2024) on ChatReport test set [^7] We fine-tune on the test instead of the training set to (1) leverage high-quality human annotation for fine-tuning; and (2) avoid indirect data leakage as $\mathcal{M}_{o}$ is fine-tuned on the training set. (fine-tuning details in App. L). We then compare embedding models’ relevance ranking with the predicted ranking of Llama3-Tok on all 43K (query, document) pairs in ClimRetrieve. We use Kendall’s $\tau$ as the metric, which directly compares the correlation between two ranks. The results are shown in Table 5. We find the Llama3-Tok-annotated benchmark successfully picks out the fine-tuned checkpoints, showing a capability of benchmarking information retrieval algorithms.
Interestingly, the unfine-tuned BGE-Base correlates more to Llama3-Tok compared to BGE-Large, although the latter shows stronger performance on MTEB (Muennighoff et al., 2023). This indicates the necessity of domain-specific benchmarking to tell the in-domain performance.

### v2

@@ 3.3 Optimizing DIRAS student LLMs
Small-embed - - - 66.34 -
Large-embed - - - 69.36 -
BGE-Gemma - - - 68.47 -
GPT-3.5 29.71 45.27 85.46 74.16 58.65
GPT-4 54.01 86.32 91.10 88.48 80.00
Llama3-CoT-Ask 36.57 76.58 89.30 86.15 72.15
Llama3-CoT-Tok 41.74 76.58 86.61 85.96 72.72
Llama3-Ask 40.18 82.11 90.14 86.02 74.61
Llama3-Tok 41.60 82.11 91.35 89.19 76.06
Phi3-CoT-Ask 36.08 72.95 88.76 80.56 69.59
Phi3-CoT-Tok 35.49 72.95 84.20 80.64 68.32
Phi3-Ask 32.30 73.23 85.56 80.05 67.79
Phi3-Tok 38.00 73.23 89.52 86.94 71.92
Gemma-CoT-Ask 31.60 72.38 86.38 81.39 67.94
Gemma-CoT-Tok 39.03 72.38 83.49 80.33 68.81
Gemma-Ask 25.74 67.13 81.80 77.43 63.02
Gemma-Tok 50.72 67.13 90.07 81.17 72.27
@@
Table 2: Comparison between the fine-tuned student $\mathcal{M}_{s}$ and different baselines on ChatReportRetrieve test data. The best scores are bolded and the second bests are underlined.$\dagger$ denotes the best score achieved by each backbone LLM.
@@
DIRAS student LLMs $\mathcal{M}_{s}$ will be used to annotate all (query, document) combinations. Two methodological choices might influence the quality of fine-tuned student LLMs. First, we explore the role of Chain-of-Thought (CoT) reasoning. Second, we investigate the choice of calibration method (Tian et al., 2023). To explore these aspects, we fine-tune $\mathcal{M}_{s}$ in four settings: $\mathcal{M}_{s}$-CoT-Ask, $\mathcal{M}_{s}$-CoT-Tok, $\mathcal{M}_{s}$-Ask, $\mathcal{M}_{s}$-Tok, where CoT means $\mathcal{M}_{s}$ is tuned to generate [Reason], [Guess], and [Confidence]; without CoT denotes $\mathcal{M}_{s}$ is tuned to only generate [Guess] and [Confidence]; “Ask” means the result is calibrated by the generated confidence score in [Confidence] field; and “Tok” means we take the token-level probability of “Yes/No” after “[Guess]:” as the confidence score for calibration. The prompt in Fig. 3 is used for fine-tuning. The “[Reason]:” line is removed in settings without CoT.
@@
We fine-tune Llama-3-8B-instruct (AI@Meta, 2024), gemma-7b-it (Team et al., 2024b), and Phi-3-mini-4k-instruct (Abdin et al., 2024) (details in App. F). We compare these fine-tuned student models with baselines including GPT-3.5 and GPT-4 using prompt $\mathcal{P}$; the OpenAI embedding models text-embedding-3-small, and text-embedding-3-large; and BGE Gemma reranker [^4] https://huggingface.co/BAAI/bge-reranker-v2-gemma, a popular LLM-based reranker for general domain.
As Fig. 4 shows, fine-tuning improves original models in all settings. Furthermore, Table 2 shows the results of all fine-tuned student models in comparison to all baselines. We observe that $\mathcal{M}_{s}$-Tok outperforms other settings for all LLM architectures. The best setting Llama3-Tok achieves GPT-4 level performance in calibration and IR on unseen questions and reports.
@@
Interestingly, we find that omitting the chain of thought usually leads to a performance increase for all three LLM architectures. CoT sometimes leads to a limited increase when asking for calibration (Ask), but constantly results in a performance drop when calibrated with token-level probability (Tok). Therefore, $\mathcal{M}_{s}$ should be fine-tuned without CoT for inference efficiency. Moreover, Tok rarely underperforms Ask, different from Tian et al. (2023)’s finding and our observations in Table 1. Thus, future work may consider probabilities of important tokens (e.g., Yes/No in our prompt template) as a promising calibration tool.
@@
4 Real-World Applications
Having established and benchmarked the design choices of DIRAS student LLMs, we now investigate their real-world application capabilities. First, we showcase how the DIRAS $\mathcal{M}_{s}$ can assist IR annotation in a real-world setting, leveraging the ClimRetrieve dataset (Schimanski et al., 2024b) (§ 4.1). Second, we show the DIRAS pipeline is also applicable for general domain QA (§ 4.2).
4.1 Applying to ClimRetrieve
ClimRetrieve (Schimanski et al., 2024b) records human analysts’ real-life workflow of reading full reports, searching for relevant information, and annotating useful information for climate-related questions with relevance scores 1-3. ClimRetrieve contains 43K (query, document) pairs (8K unique documents but each document in a report is multiplied by the amount of analyzed queries per report) out of which 595 are gold labels for relevant (query, document) pairs. Other not annotated (query, document) combinations might be either irrelevant or a part of annotation selection bias – a widely existing problem in IR annotation (Thakur et al., 2021). To succeed on this dataset, the IR model needs to (1) capture the analysts’ mental model about useful information (i.e., relevance definition), and (2) understand fine-grained differences in degree of relevance (score 1-3).
Since ClimRetrieve is still in the climate domain, we re-annotate its data with the best $\mathcal{M}_{s}$ in § 3 (Llama3-Tok), and explore whether $\mathcal{M}_{s}$’s annotations can (1) RQ1: reflect fine-grained differences in degree of relevance; (2) RQ2: be improved through refining relevance definitions; (3) RQ3: mitigate annotation selection bias for better IR recall evaluation; and (4) RQ4: benchmark and select IR algorithms.
@@ 4.1 Applying to ClimRetrieve
RQ1: Reflecting Fine-Grained Relevance Levels. We first evaluate Llama3-Tok’s annotation on 595 gold labels of ClimRetrieve to verify whether it can effectively recover analysts’ ranking for relevant content by understanding which documents are more helpful than others. Relevance definitions are drafted with GPT-4 with the same procedure as § 3. We report nDCG [^5] MAP can only measure binary relevance and since we only investigate relevant samples, it cannot be calculated. scores to measure the ranking performance on ClimRetrieve. Gold labels 1, 2, and 3 are assigned with relevance scores 1/3, 2/3, and 1. Besides OpenAI 3rd generation embedding models, we also have a random baseline where all (query, document) pairs are assigned a random relevance score between 0 and 1. The random baseline results are averaged over 5 random seeds (40 to 44). Importantly, all ClimRetrieve annotations are to some degree relevant, so improvement over the random baseline is challenging as the system needs to understand the trivial different degrees of relevance.
Table 3 presents different systems’ performance. There is a clear trend of outperformance of the fine-tuned Llama-3 models in this challenging setting.
@@
Table 4: Comparison of using the generic and the improved relevance definitions for ranking all ClimRetrieve (query, document) pairs.
RQ2: Improving Performance through Improving Definitions. So far, we used GPT-4 to draft the relevance definitions. To investigate the effect of improved definitions, we compare two setups: (1) The generic relevance definition: the definition drafted by GPT-4. (2) The improved relevance definition: The only way to improve the definition is to align it closer to ClimRetrieve annotators’ mental model of document relevance. We achieve this by adding relevant text samples to the prompt for generating the definition with GPT-4 (see App. I for details).
@@
After creating the improved definitions, we repeat predicting the relevance scores with Llama3-Tok. Since we involve examples with various relevance scores to improve relevance definitions, these definitions might not help distinguish the granular level of partial relevance. However, the improved definitions might especially help to distinguish relevant documents from irrelevant ones. Calculating the nDCG and MAP score for all 43K (query, document) pairs, we find evidence for this notion (see Table 4). Thus, the inclusion of improved definitions seems to enhance the performance (see App. J for details).
RQ3: Mitigating Annotation Selection Bias. ClimRetrieve employs a real-world analyst scenario. This entails that the human only selectively annotates documents that are likely to be relevant and assumes unannotated documents as irrelevant (see e.g., Thakur et al., 2021). Therefore, the dataset allows us to investigate our model’s capabilities to counteract biases. For this purpose, we sample 200 disagreements between Llama3-Tok’s annotation and the original ClimRetrieve labels. Then, we reannotate these samples with a human labler. To account for different confidence levels in Llama3-Tok’s prediction, we differentiate prediction with confidence higher or lower than 0.95.
As Table 5 indicates, the model can be successfully used to overturn decisions of unseen, as irrelevant assumed documents (91.30% for confidence $>0.95$). Strikingly, even samples annotated by humans, i.e., those labeled as relevant, can be overturned, though with a lower certainty. We attribute this to differences in the unknown mental model of the ClimRetrieve labeler and our explicit relevance definitions (for details, see App. L). However, for us, it is reaffirming to observe that the DIRAS model is consistent with its own definition. Thus, we conclude that DIRAS’ labeling is effective and helps to mitigate annotation selection bias.
ClimRetrieve (62.00%)
All Rel. Irr.
Conf $\leq 0.95$ 36.84 27.90 48.48
Conf $>0.95$ 85.48 78.18 91.30
Table 5: Accuracy of student model Llama3-Tok annotations that disagree with original ClimRetrieve relevance labels. All denotes all sampled disagreed labels. Rel. ( Irr.) denotes the subset where the original label is relevant (irrelevant). Conf $\leq 0.95$ ($>0.95$) denotes the subset where Llama3-Tok’s confidence is lower (higher) than 0.95. ClimRetrieve (62.12%) means that 62.12% of data are annotated with $>0.95$ confidence.
@@
Table 6: Different embedding models’ performance benchmarked by student model $\mathcal{M}_{s}$’s prediction on all 43K (query, document) pairs of ClimRetrieve. “ft” denotes the model is fine-tuned on in-domain data.
ELI5 ASQA QAMPARI RAG-Bench
N N@10 N@20 N N@10 N@20 N N@10 N@20 N N@10 N@20
GPT-4 48.43 15.97 17.81 64.62 38.82 46.32 56.21 27.54 35.34 42.91 31.84 41.35
Llama3-Tok 48.73 17.04 20.08 64.90 39.37 48.44 56.58 28.70 35.49 48.27 41.13 47.88
Table 7: Applying DIRAS to QA datasets, the IR performance of student model Llama3-Tok and GPT-4. N denotes nDCG.
ELI5 (85.05%) ASQA (66.32%) QAMPARI (72.66%) RAG-Bench (60.33%)
All Rel. Irr. All Rel. Irr. All Rel. Irr. All Rel. Irr.
Conf $\leq 0.95$ 62.35 48.48 71.15 67.27 71.43 66.67 74.11 73.33 74.23 58.41 38.10 63.04
Conf $>0.95$ 84.71 83.12 100.0 90.09 95.83 88.51 90.00 81.25 91.49 96.36 95.24 96.63
Table 8: Accuracy of student model Llama3-Tok annotations that disagree with original relevance labels. Same setup as Table 5.
@@
RQ4: Benchmarking IR. We use $\mathcal{M}_{s}$ to annotate all 43K ClimRetrieve datapoints and obtain a benchmarking dataset to select IR algorithms. This approach can be especially helpful when lacking human annotation and annotation selection bias is prevalent. Specifically, we compare the performance of embedding models before and after in-domain fine-tuning. If the $\mathcal{M}_{s}$-annotated benchmark gives higher scores to the fine-tuned checkpoints, that means it is capable of selecting a better model for this specific domain.
@@
For this experiment, we first fine-tune open-sourced embedding models bge-large-en-v1.5 and bge-base-en-v1.5 (Chen et al., 2024) on ChatReportRetrieve test set [^6] We fine-tune on the test instead of the training set to (1) leverage high-quality human annotation for fine-tuning; and (2) avoid indirect data leakage as $\mathcal{M}_{s}$ is fine-tuned on the training set. (fine-tuning details in App. K). We then compare embedding models’ relevance ranking with the predicted ranking of Llama3-Tok on all 43K (query, document) pairs in ClimRetrieve. We use Kendall’s $\tau$ as the metric, which directly compares the correlation between two ranks. The results are shown in Table 6. We find the Llama3-Tok-annotated benchmark successfully picks out the fine-tuned checkpoints, showing a capability of benchmarking information retrieval algorithms. Interestingly, the unfine-tuned BGE-Base correlates more to Llama3-Tok compared to BGE-Large, although the latter shows stronger performance on MTEB (Muennighoff et al., 2023). This indicates the necessity of domain-specific benchmarking to tell the in-domain performance.
4.2 Applying to QA Datasets
In this section, we apply the DIRAS pipeline to QA datasets that are widely used in RAG benchmarking. DIRAS addresses queries for broad information and IR recall. Thus, we include long-form QA datasets from ALCE (Gao et al., 2023), including ELI5 (Fan et al., 2019), ASQA (Stelmakh et al., 2023), and QAMPARI (Amouyal et al., 2023). We also include RAG-Bench (Fang et al., 2024) that consists of questions from TriviaQA (Joshi et al., 2017), WebQ (Berant et al., 2013), and Natural Questions (Kwiatkowski et al., 2019). RAG-Bench is chosen since it has labels for partial vs. full relevance, which is a focus of DIRAS. Importantly, the context relevance labels from ALCE and RAG-Bench are derived from reference answers to questions, using heuristics instead of manual annotation. Thus, they are to some extent noisy.
We first run the DIRAS pipeline [^7] Different from Climate change datasets, QA datasets do not require nuanced relevance definition. Thus we use a fixed relevance definition: “the document is helpful only if its content answers the query”. See full prompt in Fig. 16. on each dataset to obtain corresponding Llama3-Tok models. Then we compare them with the teacher model – GPT-4. Results in Table 7 show that Llama3-Tok outperforms GPT-4 in IR. Then for each dataset, we repeat the annotation selection bias assessment of RQ3 in § 4.1. Again, we sample 200 disagreements between Llama3-Tok annotation and the original (potentially noisy) relevance labels, and manually check whether Llama3 or original labels are correct. As shown in Table 8, Llama3-Tok’s annotations are predominately correct with a confidence $>0.95$ (e.g., $85.05\%$ of ELI5). When there is a disagreement, relying on Llama3-Tok leads to less error (Acc. $>50\%$), especially when the confidence $>0.95$, thanks to the good calibration. For ASQA, QAMPARI, and RAG-Bench, the majority ($>90\%$) of the disagreement lies in originally irrelevant labeled part of the dataset (Irr.), possibly due to (query, document) pairs are selectively annotated. DIRAS achieves high accuracy in Irr. disagreements. Therefore, we reaffirm the notion that applying DIRAS to annotate broader (query, document) pairs can effectively reduce annotation selection bias, and thus improve IR recall benchmarking. All implementation details are in App. M.